OpenStudy (anonymous):

A cereal box is 8 inches long, 3 inches wide, and 12 inches tall. (a) Volume? (b) Surface area?

OpenStudy (anonymous):

a) volume = 3*8*12 = 288 inches cubed. b) sa = 2(3*8) + 2(12*8) + 2(3*12) = 312 inches squared. hope this helps.